In the school,there are 60% more girls than boys.out of those boys,there are 50% more boys who are more right handed than left handed. If there are 2 left handed boys, how many people are in the school?
Answer by robertb(5830)

Let g = # of girls in school
b = # of boys in school,
R = # of right-handed boys in school, and L = # of left-handed boys in school.
==> g = 1.6b
Also, R = 1.5L
Now it is given that L = 2.
==> R = 1.5*2 = 3
==> number of boys is b = R + L = 3+2 = 5
==> g = 1.6*5 = 8
==> #number of boys and girls in school is b + g = 5 + 8 = 13.